OpenStudy (vanz):

find all rational roots for P(x)=0 , P(x)=2x^3-3x^2+x-1

OpenStudy (unklerhaukus):

there is only one right,

OpenStudy (anonymous):

\[2x^3-3x^2+x-1= -1x+x -1 = 2x-1 answer!!\]

OpenStudy (anonymous):

\[2x^3-3x^2\neq-1x\]

OpenStudy (anonymous):

only one root is rational : 1.398

OpenStudy (unklerhaukus):

\[\frac{\text{~seven}}{{\text{five}}}\]

OpenStudy (anonymous):

there is no minus in the answer I guess..:) It's a positive number.

OpenStudy (anonymous):

There aren't any. The rational zero theorem tells you to take the last coefficient and divide it by the first coefficient. So \[a _{n}=2\]\[a _{0}=-1\]\[a _{0}/a _{n}=\pm1/2\] So check x=1/2 and x=-1/2. \[P(1/2)\neq0\]\[P(-1/2)\neq0\]